OpenStudy (anonymous):

The following data values represent a population. What is the variance of the values? 6, 10, 14, 2 A. 20 B. 16 C. 8 D. 10

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

Can you find the expected value first?

OpenStudy (anonymous):

The mean, I mean

OpenStudy (anonymous):

how u find mean wit even set

OpenStudy (anonymous):

\[ \frac{1}{|A|}\sum_{a\in A}a \]

OpenStudy (anonymous):

o.0 wut is that

OpenStudy (anonymous):

In this case: \[ \frac{6+10+14+2}{4} \]

OpenStudy (anonymous):

pls explain

OpenStudy (anonymous):

The mean is the sum of all the numbers, divided by how many numbers there are.

OpenStudy (anonymous):

sum is add

OpenStudy (anonymous):

32/4?

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

thank u, can u help with more? im still struggling

OpenStudy (anonymous):

The mean is 8.

OpenStudy (anonymous):

Now for variance, we have to plug each value into (8-x)^2 add them, and then divide by the number of numbers. The 8 is there because it is the mean.

OpenStudy (anonymous):

The mean is \(\mu\):\[ \mu =\frac{1}{|X|}\sum_{x\in X}x \]The variance is \(\sigma^2\):\[ \sigma^2=\frac{1}{|X|}\sum_{x\in X}(x-\mu)^2 \]Where as \(\sigma\) is standard deviation.